OpenStudy (anonymous):

Find the slope of the line that passes through (-1,3) and (4,-6)

pooja195 (pooja195):

\[\LARGE \frac{y_2-y_1}{x_2-x_1}=slope\]

pooja195 (pooja195):

\[\LARGE \frac{-6-3}{-4--1}=slope\] \[\LARGE \frac{-6-3}{-4+1}=slope\]

OpenStudy (anonymous):

Thank you all

pooja195 (pooja195):

yw